Ticket Number: HEA-NOC/20060705-4 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20060705 10:57 UTC+1 Problem Start: 20060705 10:38 UTC+1
Ticket Update: 20060711 13:11 UTC+1 Problem End: 20060707 12:23 UTC+1
Site/line: Liberties VEC
Interface HEAnet-Liberties not operational
No connectivity with Liberties
20060705 10:53 LN
Liberties Serial 3/5 went down at 10:38.
I called Denis at Liberties, no response.
try to contact BT.
20060705 11:04 LN
contacted BT, they are investigating.
their ref. 1-330511926
20060705 11:18 LN
BT called their system creshed (BT), given a new ticket ref 1-330511932
Called BT again, Eircom Engineers working on it remotely.
20060705 16:04 LN
BT called back, there was a power break around the arear in this afternoon
Eircom sees no problem with the circuit.
try to contact Liberties.
20060705 17:05 DL Talk to Denis on phone and try to troubleshoot.
- Router reloaded and boots properly
- Serial transmission indicators (DCD=up DSR=up DTR=up RTS=up CTS=up)
- But carrier transitions every 10 seconds
- HDLC keepalives are sent from HEAnet but not bounced back from Liberties
debug serial interface
- HDLC keepalives are received in HEAnet from Liberties
- Could it be that HDLC link is unidirectional?
HEAnet -> Liberties does not work
Liberties -> HEAnet does work
20060706 10:08 DW BT phone to report that their system came back up and
the fault is now logged with Eircom for a line check.
20060706 12:14 LN BT called back, Eircom see no error on that circuit.
20060706 14:50 DL Requested support from Lancomms and router replacement.
Communicate it to Denis McCarthy
20060707 12:20 LN
BT called back The problem has been fixed, connectivity has regained.
20060707 20:02 COC Liberties have loss connectivity again 17:49.
20060707 20:18 COC Circuit is flapping. Contacted BT who are testing the
line between 20:30 and 22:30
20060708 12:24 COC Circuit was tested between 20060707 22:30 and 00:30.
The line passed the required tests. This link is still transitioning.
Contacted BT again.
20060708 17:03 COC Circuit is being tested again from 17:00 - 19:00 on the
20060709 12:43 COC No faults have been found with the serial line after 4
hour of testing. Contacted Lancomms to send replacement parts
20060710 13:09 COC Lancomms have replaced both router and serial cable but
errors still were still being logged on both interfaces of the circuit.
20060711 13:08 COC BERT test scheduled for 20060711 14:00 - 15:00. There
will be no connectivity to Liberties during this period.
This ticket can be monitored at http://www.hea.net/tickets/20060705-4